OpenStudy (anonymous):

evaluate the function for x=-2a, if a=3 t(x)=-4x+7 a.31 b.-4 c.99 d.-17

OpenStudy (anonymous):

well first find x,what is x?

OpenStudy (anonymous):

x = -2a = -2(3) = ?

OpenStudy (anonymous):

I really don't know

OpenStudy (anonymous):

u don't know what is (-2)(3) = ?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

i just need the answer

OpenStudy (anonymous):

no.don't just ask for answer! OpenStudy is to learn!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

and -2(3) = -6 so x = -6 and for x we have f(x) = -4x + 7 so find f(x) that x = -6 (you should substitute -6 instead x here: -4x +7

OpenStudy (anonymous):

so the answer is 31

OpenStudy (anonymous):

yes ;) correct

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

you're welcome ;) you can click on best response button to give me a medal and close the topic as u got the answer. :)